What percent of babies weigh over 3600 grams?


1
Expert's answer
2020-10-02T09:26:18-0400

Babies weight is the normal distribution. Mean value - 3400 grams, standard deviation - 550 grams.

P(x>3600)=1-P(x<3600)

P(x<3600)

z=(x-mean)/sd=(3600-3400)/550=0.36, using table, P(x<3600) = 0.64,so

P(x>3600)=1-0.64=0.36=36%